OpenStudy (anonymous):

horizontal shift of: -240sin(4θ-360) and 4cos(1/2θ-15)-2.5 Please explain how you got those answers! Thanks :)

OpenStudy (paxpolaris):

If original is y= f(x) and final is, \[y-k = a \cdot f(x-h)\]or y = a⋅f(x−h) + k h is horizontal shift k is vertical shift a is stretch/ squish

OpenStudy (anonymous):

So in this case, -360 would be considered the horizontal shift, and it doesn't consist of a vertical shift?

OpenStudy (paxpolaris):

+360 is horizontal shift from -240sin(4θ)

OpenStudy (anonymous):

Why is it positive?

OpenStudy (paxpolaris):

if f(x) becomes f(x-h) because now f(1) is moved to f(1+h) ... anyways if θ is in degrees it does not matter because -240sin(4θ-360) will look exactly like -240sin(4θ)

OpenStudy (anonymous):

So basically what ever h's sign is in the brackets, should be reversed? ex: -3cos(θ-15) would be positive 15?

OpenStudy (paxpolaris):

yep

OpenStudy (anonymous):

Ahhh! Okay Thanks a lot!